# UniST: A Prompt-Empowered Universal Model for Urban Spatio-Temporal Prediction

### v3

@@ 5.5. Zero-Shot Prediction
Results.
Figure 4 also compare the performance of UniST (zero-shot) and baselines (few-shot). As observed, UniST achieves remarkable zero-shot performance, even surpassing many baselines trained with training data that are highlighted by red dashed lines. We attribute these surprising results to the powerful spatio-temporal transfer capability. It suggests that for a completely new scenario, even when the displayed overall patterns are dissimilar to the data encountered during the training process, UniST can extract fine-grained similar patterns from our defined spatial and temporal properties. The few-shot and zero-shot results demonstrate the data-efficient advantage of UniST.
6. Analysis
@@ 6.1. Ablation Study
The prompts play an essential role in our UniST model. Here we investigate whether the designed spatial and temporal properties contribute to the final performance. We use ’s’ to denote spatial closeness and hierarchy, ’p’ for temporal periodicity, and ’c’ for temporal closeness. we compare the overall design that incorporates all three properties with three degraded versions that individually remove ’s’, ’p’, or ’c’. Figure 5(a) shows the results on four traffic speed datasets. As we can observe, removing any property results in a performance decrease. The contributions of each spatial and temporal property vary across different datasets, highlighting the necessity of each property for the spatio-temporal design.
Additionally, we explore how the number of embeddings in the memory pools affects the final performance. As seen in Figure 5(b), increasing the number from 128 to 512 improves performance across the four datasets. When further increasing the number to 1024, the performance remains similar to 512, suggesting that 512 is an optimal choice.
Figure 6. Visualization of embeddings in the spatial and temporal memory pools at the initial and final optimized states. The embeddings exhibit obvious divergence.
Figure 7. (a) and (b): Comparison of the mean value of inputs in each memory embedding, where the inputs assign the highest attention weight to the memory embedding. Different cities exhibit similar distributions. (c) and (d): Comparison of the attention weight on each memory embedding for two distinct datasets. The two datasets showcase remarkably different distributions.
6.2. Analysis of Prompt Learner
@@
Figure 8. Left: training loss across five models with varying parameter sizes. Right: performance evaluation of masked patch reconstruction by increasing parameter sizes.
@@ 6.3. Scalability
We analyze the scalability of our UniST model. Our investigation specifically concentrates on observing changes in training loss and prediction performance as we vary the model parameter size. Figure 8 depicts the training loss and testing RMSE of UniST with varying parameter sizes. For training loss (left figure), we observe two key characteristics: (i) across different parameter sizes, the training loss consistently decreases and gradually converges with increasing training steps; (ii) increazing the parameter size accelerates the convergence of the training loss; (iii) there exist diminishing marginal returns, indicating that the training loss cannot be indefinitely reduced. The right figure illustrates the reconstruction RMSE on the testing set, showing similar trends to the training loss.
@@ 7. Conclusion
In this work, we address an important problem: building a universal model UniST for urban spatio-temporal prediction. By leveraging the diversity of spatio-temporal data from multiple sources, and discerning and aligning underlying shared commonalities of spatio-temporal patterns across multiple scenarios, UniST demonstrates a powerful capability to predict across all scenarios, particularly in few-shot and zero-shot settings. The limitation of UniST lies in its reliance on grid-based spatial partitioning. A promising direction for future work entails the integration of various spatio-temporal data formats, such as grid-based, sequence-based, and graph-based data. In essence, the universality of deep learning models holds the promise to bridge the gap between artificial and human intelligence. We believe our study will inspire future research in spatio-temporal modeling towards the universal direction.
@@ Appendix A Datasets
Here we provide more details of the used datasets in our study. We collect various spatio-temporal data from multiple cities and domains. Table 3 summarizes the basic information of the used datasets, and Table 4 reports the basic statistics. We have normalized all datasets to the range $[-1,1]$. The reported prediction results are denormalized results. Specifically, values for Crowd and Cellular datasets in Table 1, Table 2, Table 8, Table 9 and Figure 4 should be scaled by a factor of $10^{3}$.
@@ Appendix B Algorithms
Appendix C Baselines
  • MC-STL (Zhang et al., 2023b): It leverages an state-of-the-art training techniques for spatio-temporal predition, the mask-enhanced contrastive learning, which can effectively capture the relationships on the spatio-temporal dimension.
  • SimVP (Gao et al., 2022a): It is a simple yet very effective video predcition model. It is completely built based on convulutional neural networks and uses MSE loss. It serves as a solid baseline in video prediction tasks.
  • STID (Shao et al., 2022): It is a MLP-based spatio-temporal prediction model, which is simple yet effective. Its superior performance comes from the identification of the indistinguishability of samples in spatio-temporal dimensions. It demonstrates that it is promising to design efficient and effective models in spatio-temporal predictions.
  • iTransformer (Liu et al., 2023b): This is the state-of-the-art multivariate time series model. Different from other Transformer-based methods, it employes the attention and feed-forward operation on an inverted dimension, that is, the multivariate correlation.
@@ Appendix E Additional Results
TaxiNYC-1 BikeNYC2 TaxiNYC-2 TrafficBJ TrafficNJ TrafficWH TrafficSZ
@@
Table 9. Performance comparison in few-shot and zero-shot (only UniST) learning settings on the Crowd dataset in terms of MAE and RMSE. 1% , 5%, and 10% denote that only the percentage of training data is utilized. We use the average prediction errors over all prediction steps.
@@
Table 10. Performance comparison in few-shot and zero-shot (only UniST) learning settings on the BikeNYC dataset in terms of MAE and RMSE. 1% , 5%, and 10% denote that only the percentage of training data is utilized. We use the average prediction errors over all prediction steps.
@@
Table 11. Performance comparison in few-shot and zero-shot (only UniST) learning settings on the TaxiBJ dataset in terms of MAE and RMSE. 1% , 5%, and 10% denote that only the percentage of training data is utilized. We use the average prediction errors over all prediction steps.

### v4

@@ 5.1. Ablation Study
The prompts play an essential role in our UniST model. Here we investigate whether the designed spatial and temporal properties contribute to the overall performance. We use ‘s’ to denote spatial closeness and hierarchy, ‘p’ for temporal periodicity, and ‘c’ for temporal closeness. we compare the overall design that incorporates all three properties with three degraded versions that individually remove ‘s’, ‘p’, or ‘c’. Figure 5(a) shows the results on four traffic speed datasets. As we can observe, removing any property results in a performance decrease. The contributions of each spatial and temporal property vary across different datasets, highlighting the necessity of each property for the spatio-temporal design.
Additionally, we explore how the number of embeddings in the memory pools affects the final performance. As seen in Figure 5(b), increasing the number from 128 to 512 improves performance across the four datasets. When further increasing the number to 1024, the performance remains similar to 512, suggesting that 512 is the optimal choice.
Figure 6. Embeddings visualization of spatial and temporal memory pools at the initial and final optimized states. The embeddings exhibit obvious divergence.
Figure 7. (a) and (b): Comparison of the mean value of inputs in each memory embedding, where the inputs assign the highest attention weight to the memory embedding. (c) and (d): Comparison of the attention weight on each memory embedding for two distinct datasets.
5.2. Prompt Learner
@@
Figure 8. (a) Training loss across five models with varying parameter sizes. (b) Performance evaluation of masked patch reconstruction by increasing parameter sizes.
@@ 5.3. Scalability
Scalability is a crucial characteristic for universal models, therefore, we explore the scaling behavior of our UniST model. Our investigation specifically concentrates on observing changes in training loss and prediction performance as we vary the model parameter size. Figure 8 depicts the training loss and testing RMSE of UniST with varying parameter sizes. Regarding training loss (left figure), several key observations emerge: (i) across different parameter sizes, the training loss consistently decreases and gradually converges with increasing training steps; (ii) increasing the parameter size accelerates the convergence of the training loss; (iii) there exist diminishing marginal returns, suggesting that reducing the training loss becomes progressively harder as parameter size increases. The right figure illustrates the reconstruction RMSE on the testing set, showing similar trends to the training loss.
These observations indicate that UniST has shown scalability behaviors,wherein larger models generally exhibit improved performance. However, unlike large language and vision models (Kaplan et al., 2020; Bai et al., 2023), the scalability in spatio-temporal prediction shows diminishing marginal returns. This may stem from the relative lack of diversity in spatio-temporal data compared to language or visual datasets.
@@ 6. Conclusion
In this work, we address an important problem of building a universal model UniST for urban spatio-temporal prediction. By leveraging the diversity of spatio-temporal data from multiple sources, and discerning and aligning underlying shared spatio-temporal patterns across multiple scenarios, UniST demonstrates a powerful capability to predict across all scenarios, particularly in few-shot and zero-shot settings. A promising direction for future work entails the integration of various spatio-temporal data formats, such as grid, sequence, and graph data. Our study inspires future research in spatio-temporal modeling towards the universal direction.
Acknowledgments
This work was supported in part by the National Key Research and Development Program of China under grant 2020YFA0711403 and the National Natural Science Foundation of China under 62171260 and 62272260.
@@ Appendix A Datasets
A.1. Basic Information
Here we provide more details of the used datasets in our study. We collect various spatio-temporal data from multiple cities and domains. Table 4 summarizes the basic information of the used datasets, and Table 15 (arXiv) reports the basic statistics. Specifically, values for Crowd and Cellular datasets in Table 2, Table 3, Table 13 (arXiv), Table 14 (arXiv) and Figure 4 should be scaled by a factor of $10^{3}$.
A.2. Data Preprocessing
For each dataset, We split it into three non-overlapping periods: the first 70% of the period was used as the training set, the next 15% as the validation set, and the final 15% as the test set. To ensure no overlap between train/val/test sets, we removed intermediate sequences. We have normalized all datasets to the range $[-1,1]$. The reported prediction results are denormalized results.
@@ A.2. Data Preprocessing
Appendix B Baselines
  • MC-STL (Zhang et al., 2023a): It leverages an state-of-the-art training techniques for spatio-temporal predition, the mask-enhanced contrastive learning, which can effectively capture the relationships on the spatio-temporal dimension.
  • SimVP (Gao et al., 2022): It is a simple yet very effective video predcition model. It is completely built based on convulutional neural networks and uses MSE loss. It serves as a solid baseline in video prediction tasks.
  • STID (Shao et al., 2022b): It is a MLP-based spatio-temporal prediction model, which is simple yet effective. Its superior performance comes from the identification of the indistinguishability of samples in spatio-temporal dimensions. It demonstrates that it is promising to design efficient and effective models in spatio-temporal predictions.
  • iTransformer (Liu et al., 2023a): This is the state-of-the-art multivariate time series model. Different from other Transformer-based methods, it employes the attention and feed-forward operation on an inverted dimension, that is, the multivariate correlation.
  • PromptST (Zhang et al., 2023c): It is the state-of-the-art pre-trianing and prompt-tuning approach for spatio-temporal prediction.
@@ D.2. Parameter Settings
D.3. Prompt-Tuning
The prompt-tuning stage aims to train a effective prompt network, which generates customized prompt for specific spatio-temporal pattern. We propose to leverage four types of spatio-temporal knowledge: (i) spatial closess ($s_{c}$), (ii) spatial hierarchy ($s_{h}$), (iii) temporal closeness ($t_{c}$), and (iv) temporal period ($t_{p}$). These knowledge-guided features are extracted from the input sequence. The input is the historical spatio-temporal sequence, the output is the predicted future spatio-temporal sequence, and the objective is to minimize the distance between the predicted results and real data. Specifically, we use the widely adopted mean squared error loss function with $l_{2}$ regularization on the parameters in UniST to prevent over-fitting, which can be formulated as follows
(2) $\mathcal{L}=\frac{1}{M}\sum(\hat{y}-y)^{2}+\lambda\sum_{\theta\in\Theta}\|\mathcal{\theta}\|_{2}$
where $\hat{y}$ and $y$ are ground truths and model predictions, respectively; $\Theta$ denotes the set that contains all model parameters.
D.4. Baseline Implementation
We compare UniST with a broad collection of state-of-the-art models for spatio-temporal prediction, which can be categorized into five groups as introduced in Section 4.1. If we consider the scalability to diverse data formats, i.e., different spatio-temporal data shapes, these baselines can be categorized into two groups: (i) approaches that are scalable with different spatio-temporal scales, such as PatchTST, MAML, and MetaST, and (ii) approaches that are non-scalable, including deep urban prediction approaches, video prediction approaches, and iTransformer. Most baselines are not scalable to different data shapes because they require a fixed number of spatial grids or variables, as seen in CNN-based approaches, MLP-based approaches, and multivariate time series models. Due to the varied data shapes, non-scalable baselines cannot be trained using all datasets, so we train separate models for each dataset.
For the scalable baseline, PatchTST (Nie et al., 2022), it utilizes a channel-independent patch time series Transformer architecture, allowing it to be applied to datasets with varied spatio-temporal shapes. To ensure a fair comparison, we train both separate models and a single ”one-for-all” model, as shown in Table 2.
Notably, there are two baselines employ pretraining and finetuning: PatchTST (Nie et al., 2022) and PromptST (Zhang et al., 2023c). However, PromptST requires a fixed number of nodes, limiting its flexibility across different data formats. In contrast, the channel-independence of PatchTST allows it to handle varied data shapes. While PromptST is a state-of-the-art pre-training and prompt-tuning approach, it lacks generalization ability across different datasets.
Figure 9. Comparison of attention weight distribution between the training set and testing set of the CrowdNJ dataset. The generated prompts assign attention weights on embeddings in the memory pool.
Figure 10. Comparison of attention weight distribution between the the TaxiBJ dataset and SpeedSZ dataset. The generated prompts assign attention weights on embeddings in the memory pool.
D.5. Experimental Design
In our experimental design, we incorporate four distinct prediction tasks: short-term prediction, long-term prediction, few-shot prediction, and zero-shot prediction. This design aligns with established practices in foundation models for time series forecasting (Zhou et al., 2023b; Jin et al., 2023a; Nie et al., 2022). The short-term and long-term prediction tasks are conducted without transfer learning settings. In these tasks, the model is trained on a set of $N$ datasets and then evaluated on the corresponding testing sets from these datasets. This setup enables us to directly assess the model’s performance across multiple datasets using a single universal model.
Furthermore, the few-shot and zero-shot prediction tasks are designed to evaluate the model’s generalization capabilities. In these tasks, the model learns from a set of source datasets to build a pretrained model and a memory pool, which is then utilized for prediction on target datasets. The key difference between the few-shot and zero-shot settings lies in the fine-tuning process on the target dataset. In few-shot prediction, the model undergoes a limited fine-tuning process using a small percentage of the target dataset’s training data, while in zero-shot prediction, the model directly applies the pre-trained model and memory pool to make predictions on the target dataset without any fine-tuning.
These four tasks collectively offer a comprehensive evaluation of the model’s performance and its ability to generalize across diverse spatio-temporal datasets.
Figure 11. Comparison of attention weight distribution between the the CrowdNJ dataset and the CellularNJ dataset. The generated prompts assign attention weights on embeddings in the memory pool.
@@ Appendix E Additional Results
E.1. Analysis of Distribution Shifts
Here, we delve into a detailed analysis of the generated prompts across different datasets. For each dataset, we compute the attention weights on the embeddings in the memory pool and visualize the distribution of these weights in Figure 9 to Figure 10. We have selected three typical scenarios to explore:
  1. (1)
    Training and Testing Sets of One Dataset: This analysis aims to investigate the model’s ability to generalize within a familiar dataset.
  2. (2)
    Two Datasets from Different Domains in the Same City: Understanding how the model adapts its prompt generation across different but related datasets can provide insights into its domain-specific learning.
  3. (3)
    Datasets from Different Cities and Domains: This scenario highlights the model’s ability to leverage knowledge learned previously and generate useful prompts adaptively.
As shown in Figure 9 to Figure 10, our analysis reveals compelling insights into the effectiveness of our prompting mechanism in handling distribution shifts. Specifically, we observed that similar prompts are consistently generated for datasets exhibiting similar spatio-temporal patterns. For instance, the prompts generated for the training and testing sets of a single dataset, as well as for the testing sets of two datasets from different domains within the same city, are similar. This consistency in prompt generation suggests that our model effectively captures and leverages the underlying spatio-temporal patterns shared between these datasets. Meanwhile, our model generates distinct prompts for scenarios involving datasets from different cities and domains, indicating its ability to adapt to diverse spatio-temporal contexts. This adaptability is crucial for handling distribution shifts, as it allows the model to flexibly adjust its prompt generation strategy based on the unique characteristics of each dataset.
Table 7. Performance on different noise levels with sine-cosine positional encoding.
TaxiBJ Crowd Cellular BikeNYC TrafficSH
Noise level
0 26.841 3.00 14.294 3.506 0.6650
0.1% 26.846 3.038 14.297 3.507 0.6651
1% 26.90 3.039 14.390 3.534 0.6653
10% 28.76 3.29 14.91 3.695 0.6877
Best baseline 27.36 3.85 16.48 3.93 0.742
Table 8. Performance on different noise levels with learnable positional encoding.
TaxiBJ Crowd Cellular BikeNYC TrafficSH
Noise level
0 27.02 3.31 15.054 3.609 0.686
0.1% 27.032 3.310 15.068 3.607 0.6860
1% 27.29 3.589 16.544 3.696 0.6911
10% 43.80 11.436 70.360 8.173 1.228
Best baseline 27.36 3.85 16.48 3.93 0.742
Table 9. Comparison of computational cost and memory cost between different approaches. The training time denotes the time cost to train all instances with one epoch.
Model STResNet ACFM STID STNorm STGSP MC-STL MAU PredRNN MIM SimVP TAU PatchTST iTransformer UniST
Model Size (M) 2.51 1.90 1.63 1.15 5.51 6.35 10.55 17.07 26.24 9.96 9.55 2.59 25.27 6.71
Memory Cost (MB) 1475 1671 1715 2539 1459 1607 1579 1065 1241 1039 1075 2859 2935 2875
Training Time (min) 0.057 0.561 0.054 0.461 0.078 0.311 0.828 0.455 0.836 0.224 0.224 0.338 0.093 1.4 (20+ datasets)
Inference Time (min) 0.011 0.026 0.007 0.070 0.006 0.013 0.026 0.015 0.024 0.013 0.010 0.031 0.012 0.034
Figure 12. Visualization of different spatio-temporal datasets: Firstly, the high-dimensional data is reduced to a two-dimensional vector using t-SNE. Subsequently, the embeddings are visualized in clusters using the k-means clustering method.
E.2. Performance under Noise Perturbations
The model’s ability to handle noisy data is necessary to ensure reliable predictions. Therefore, we conduct experiments to evaluate UniST’s robustness against noisy data. Specifically, we introduced Gaussian noise with varying levels of intensity to the input data and assessed UniST’s performance under these conditions. We considered three levels of noise: Gaussian noise randomly sampled from a 0.1% normal distribution, Gaussian noise randomly sampled from a 1% normal distribution, and Gaussian noise randomly sampled from a 10% normal distribution. These noise levels represent varying degrees of data corruption, simulating real-world scenarios where data can be noisy or contain irregularities.
The results, as detailed in Table 7, demonstrate that UniST consistently outperforms baseline models even in the presence of noise perturbations (where the best baseline has no noise perturbation). This suggests that UniST is capable of effectively handling noisy data, which is crucial for ensuring reliable predictions, especially in real-world scenarios where data can be messy or contain irregularities.
Moreover, we examine how different positional encoding methods affect the model’s robustness. We compare the use of two positional encoding methods: learnable embeddings and sine-cosine encoding. The results in Table 8 show the performance with learnable embeddings, while Table 7 shows the performance with sine-cosine encoding. Comparing these two sets of results, we observe that sine-cosine encoding exhibits more robust performance against noise perturbations. Specifically, learnable embeddings show a significant performance reduction with increased noise perturbation and perform worse than the best baseline model.
E.3. Model Efficiency
Table 9 shows a detailed comparison of the computational and memory costs of UniST against baselines. The results show that the model size and memory cost of UniST are comparable to those of other approaches. However, due to the multiple data pre-training involved, the training time of UniST is longer compared to other methods. Despite this, UniST consistently outperforms baselines on all datasets with just one model. Thus, we consider the additional training time acceptable given the superior performance achieved.
E.4. Dataset Similarity
To assess the similarities among the datasets used in our study, we employed a two-step process. First, we reduced the dimensionality of the spatio-temporal data using t-SNE, a technique for dimensionality reduction. This allowed us to visualize the datasets in a lower-dimensional space. Second, we applied the k-means clustering method to the reduced data to identify clusters of similar spatio-temporal patterns.
Figure 13. Ablation studies on four types spatial and temporal knowledge extraction $t_{c},t_{p},s_{c}$, and $s_{h}$.
Table 10. Ablation studies on four masking strategies.
Prediction Imputation Spatial extrapolation
Complete 0.781 0.761 0.729
wo/ Random masking 0.796 1.72 0.761
wo/ Tube masking 0.787 0.788 0.817
wo/ Block masking 0.785 0.773 1.02
wo/ Temporal masking 1.44 0.772 0.742
The results of our visualization revealed interesting insights. We found that certain datasets, such as the Crowd data and Cellular data in Nanjing, exhibited similar spatio-temporal patterns. Similarly, the Bike data and Taxi data in New York City showed similarities in their patterns. However, most datasets from different cities or domains exhibited distinct spatio-temporal patterns, indicating significant distribution shifts. These observations highlight the powerful generalization ability and universality of our approach across datasets with significantly distinct spatio-temporal patterns.
@@ E.4. Dataset Similarity
TaxiNYC-1 BikeNYC-2 TaxiNYC-2 TrafficBJ TrafficNJ TrafficWH TrafficSZ
@@
E.5. Additional Ablation Studies
E.5.1. Masking Strategies.
We investigated the contribution of each of the four masking strategies by comparing the performance when all four strategies are employed with the performance when one of the strategies is removed. We conducted experiments on three spatio-temporal tasks: prediction, imputation, and spatial extrapolation, using the TrafficCD dataset.
The results, shown in Table 10, indicate that training with all four masking strategies achieved the best performance across all three tasks. Removing the temporal masking strategy results in the most significant performance decrease for the prediction task, removing the random masking strategy leads to the most significant performance decrease for the imputation task, and removing the block masking strategy results in the most significant performance decrease for the spatial extrapolation task. These results are reasonable as each masking strategy is designed to align with a specific task objective.
It is worth noting that despite the seemingly mismatched nature of some masking strategies with certain spatio-temporal tasks (e.g., random masking vs. prediction, temporal masking vs. imputation, and temporal masking vs. spatial extrapolation), we find that these masking strategies still contribute to the performance of less related tasks. This indicates that the masking strategies not only benefit their intended tasks but also have broader effects on the model’s general learning of spatio-temporal dependencies and dynamics. For example, while random masking may seem unrelated to causal prediction tasks, it can help the model learn robust features that generalize well across different time points. Additionally, temporal masking can help the model better understand the temporal dynamics when performing spatial extrapolation.
E.5.2. Knowledge-Guide Prompts.
The prompts play an essential role in our UniST model. Here we investigate whether the designed spatial and temporal properties $s_{c},s_{h},t_{c}$, and $t_{p}$ contribute to the final performance. We use $s_{c}$ to denote spatial closeness, $s_{h}$ to denote spatial hierarchy, $t_{p}$ for temporal periodicity, and $t_{c}$ for temporal closeness.
we compare the overall design that incorporates all three properties with four degraded versions that individually remove $s_{c},s_{h},t_{c}$, or $t_{p}$ Figure 13 shows the results on four traffic speed datasets. As we can observe, removing any property results in a performance decrease. The contributions of each spatial and temporal property vary across different datasets, highlighting the necessity of each property for the spatio-temporal design.
E.6. Additional Prediction Results
[…]
@@ E.6. Additional Prediction Results
Table 14. Performance comparison in few-shot and zero-shot (only UniST) settings on the Crowd dataset in terms of MAE and RMSE. 1% , 5%, and 10% denote that only the percentage of training data is utilized. We use the average prediction errors over all prediction steps.
@@
Table 15. Performance comparison in few-shot and zero-shot (only UniST) settings on the BikeNYC dataset in terms of MAE and RMSE. 1% , 5%, and 10% denote that only the percentage of training data is utilized. We use the average prediction errors over all prediction steps.
@@
Table 16. Performance comparison in few-shot and zero-shot (only UniST) settings on the TaxiBJ dataset in terms of MAE and RMSE. 1% , 5%, and 10% denote that only the percentage of training data is utilized. We use the average prediction errors over all prediction steps.